# FuseMoE: Mixture-of-Experts Transformers for Fleximodal Fusion

### v2

@@ I.3 Over-specified Setting
Arguing in a similar fashion as in Step 3 of Appendix K.2, we obtain that $\xi_{j,\alpha_{1},\alpha_{2},\eta}=\kappa_{j,\gamma}=0$ for any $j\in[K]$, $0\leq|\alpha_{1}|+|\alpha_{2}|\leq 2\bar{r}_{j}$, $0\leq\eta\leq 2(\bar{r}_{j}-|\alpha_{1}|-|\alpha_{2}|)$ and $0\leq|\gamma|\leq 1+\mathbf{1}_{\{|\mathcal{A}_{j}|>1\}}$. This contradicts the fact that at least one among them is non-zero. Hence, the proof is completed.
Appendix L Auxiliary Results
Lemma L.1.
For any $i\in[k_{*}]$, let $W_{i},W^{*}_{i}\in\mathbb{R}^{d}$ such that $\|W_{i}-W^{*}_{i}\|\leq\eta_{i}$ for some sufficiently small $\eta_{i}>0$. Then, for any $\tau\in[q]$, unless the set $\mathcal{X}^{*}_{\tau}$ has measure zero, we obtain that $\mathcal{X}^{*}_{\tau}=\mathcal{X}_{\tau}$ where
$\mathcal{X}_{\tau}:=\{x\in\mathcal{X}:-\|W_{i}-x\|\geq-\|W_{i^{\prime}}-x\|,\forall i\in\{\tau_{1},\ldots,\tau_{K}\},i^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}\}.$
Proof of Lemma L.1.
Let $\eta_{i}=M_{i}\varepsilon$, where $\varepsilon$ is some fixed positive constant and $M_{i}$ will be chosen later. For an arbitrary $\tau\in[q]$, since $\mathcal{X}$ and $\Omega$ are bounded sets, there exists some constant $c^{*}_{\tau}\geq 0$ such that
$\min_{x,i,i^{\prime}}\Big{[}-\|W^{*}_{i}-x\|+\|W^{*}_{i^{\prime}}-x\|\Big{]}=c^{*}_{\tau}\varepsilon,$ (31)
where the minimum is subject to $x\in\mathcal{X}^{*}_{\tau},i\in\{\tau_{1},\ldots,\tau_{K}\}$ and $i^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}$. We will point out that $c^{*}_{\tau}>0$. Assume by contrary that $c^{*}_{\tau}=0$. For $x\in\mathcal{X}^{*}_{\tau}$, we may assume for any $1\leq i<j\leq k_{*}$ that
$-\|W^{*}_{\tau_{i}}-x\|\geq-\|W^{*}_{\tau_{j}}-x\|.$
Since $c^{*}_{\tau}=0$, it follows from equation 31 that $\|W^{*}_{\tau_{K}}-x\|=\|W^{*}_{\tau_{K+1}}-x\|$. In other words, $\mathcal{X}^{*}_{\tau}$ is a subset of $\mathcal{Z}:=\{x\in\mathcal{X}:\|W^{*}_{\tau_{K}}-x\|=\|W^{*}_{\tau_{K+1}}-x\|\}$. As $W^{*}_{\tau_{K}}\neq W^{*}_{\tau_{K+1}}$ and the distribution of $X$ is continuous, it follows that the set $\mathcal{Z}$ has measure zero. Since $\mathcal{X}^{*}_{\tau}\subseteq\mathcal{Z}$, we can conclude that $\mathcal{X}^{*}_{\tau}$ also has measure zero, which contradicts the hypothesis of Lemma L.1. Thus, we must have $c^{*}_{\tau}>0$.
As $\mathcal{X}$ is a bounded set, we assume that $\|x\|\leq B$ for any $x\in\mathcal{X}$. Let $x\in\mathcal{X}^{*}_{\tau}$, then we have for any $i\in\{\tau_{1},\ldots,\tau_{K}\}$ and $i^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}$ that
$-\|W_{i}-x\|$ $=-\|W_{i}-W^{*}_{i}\|-\|W^{*}_{i}-x\|$
$\geq-M_{i}\varepsilon-\|W^{*}_{i^{\prime}}-x\|+c^{*}_{\tau}\varepsilon$
$=-M_{i}\varepsilon+c^{*}_{\tau}\varepsilon-\|W^{*}_{i^{\prime}}-W_{i^{\prime}}\|-\|W_{i^{\prime}}-x\|$
$\geq-2M_{i}\varepsilon+c^{*}_{\tau}\varepsilon-\|W_{i^{\prime}}-x\|.$
By setting $M_{i}\leq\dfrac{c^{*}_{\tau}}{2}$, we get that $x\in{\mathcal{X}}_{\tau}$, which means that $\mathcal{X}^{*}_{\tau}\subseteq{\mathcal{X}}_{\tau}$. Similarly, assume that there exists some constant $c_{\tau}\geq 0$ that satisfies
$\min_{x,i,i^{\prime}}\Big{[}-\|W_{i}-x\|+\|W_{i^{\prime}}-x\|\Big{]}=c^{*}_{\tau}\varepsilon.$
Here, the above minimum is subject to $x\in\mathcal{X}_{\tau}$, $i\in\{\tau_{1},\ldots,\tau_{K}\}$ and $i^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}$. If $M_{i}\leq\dfrac{c_{\tau}}{2}$, then we also receive that $\mathcal{X}_{\tau}\subseteq\mathcal{X}^{*}_{\tau}$. Hence, if we set $M_{i}=\dfrac{1}{2}\min\{c^{*}_{\tau},c_{\tau}\}$, we reach the conclusion that $\mathcal{X}^{*}_{\tau}=\mathcal{X}_{\tau}$. ∎
Lemma L.2.
For any $j\in[k_{*}]$ and $i\in\mathcal{A}_{j}$, let $W_{i},W^{*}_{j}\in\mathbb{R}^{d}$ that satisfy $\|W_{i}-W^{*}_{j}\|\leq\eta_{j}$ for some sufficiently small $\eta_{j}>0$. Additionally, for $\max_{\{\tau_{j}\}_{j=1}^{K}\subseteq[k_{*}]}\sum_{j=1}^{K}|\mathcal{A}_{\tau_{j}}|\leq\overline{K}\leq k$, we assume that there exist $\tau\in[q]$ and $\overline{\tau}\in[\overline{q}]$ such that $\{\overline{\tau}_{1},\ldots,\overline{\tau}_{\overline{K}}\}=\mathcal{A}_{\tau_{1}}\cup\ldots\cup\mathcal{A}_{\tau_{K}}$. Then, if the set $\mathcal{X}^{*}_{\tau}$ does not have measure zero, we achieve that $\mathcal{X}^{*}_{\tau}=\overline{\mathcal{X}}_{\overline{\tau}}$.
Proof of Lemma L.2.
Let $\eta_{j}=M_{j}\varepsilon$, where $\varepsilon$ is some fixed positive constant and $M_{i}$ will be chosen later. As $\mathcal{X}$ and $\Omega$ are bounded sets, we can find some constant $c^{*}_{\tau}\geq 0$ such that
$\min_{x,j,j^{\prime}}\Big{[}-\|W^{*}_{j}-x\|+\|W^{*}_{j^{\prime}}-x\|\Big{]}=c^{*}_{\tau}\varepsilon,$
where the above minimum is subject to $x\in\mathcal{X}^{*}_{\tau},j\in\{\tau_{1},\ldots,\tau_{K}\}$ and $j^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}$. By arguing similarly to the proof of Lemma L.1, we deduce that $c^{*}_{\tau}>0$.
Let $x\in\mathcal{X}^{*}_{\tau}$ and $\overline{\tau}\in[\overline{q}]$ such that $\{\overline{\tau}_{1},\ldots,\overline{\tau}_{\overline{K}}\}=\mathcal{A}_{\tau_{1}}\cup\ldots\cup\mathcal{A}_{\tau_{K}}$. Then, for any $i\in\{\overline{\tau}_{1},\ldots,\overline{\tau}_{\overline{K}}\}$ and $i^{\prime}\in\{\overline{\tau}_{\overline{K}+1},\ldots,\overline{\tau}_{k}\}$, we have that
$-\|W_{i}-x\|$ $=-\|W_{i}-W^{*}_{j}\|-\|W^{*}_{j}-x\|$
$\geq-M_{i}\varepsilon-\|W^{*}_{j^{\prime}}-x\|+c^{*}_{\tau}\varepsilon$
$=-M_{i}\varepsilon+c^{*}_{\tau}\varepsilon-\|W^{*}_{j^{\prime}}-W_{j^{\prime}}\|-\|W_{i^{\prime}}-x\|$
$\geq-2M_{i}\varepsilon+c^{*}_{\tau}\varepsilon-\|W_{i^{\prime}}-x\|,$
where $j\in\{\tau_{1},\ldots,\tau_{K}\}$ and $j^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}$ such that $i\in\mathcal{A}_{j}$ and $i^{\prime}\in\mathcal{A}_{j^{\prime}}$. If $M_{j}\leq\dfrac{c^{*}_{\tau}}{2}$, then we get that $x\in{\mathcal{X}}_{\overline{\tau}}$, which leads to $\mathcal{X}^{*}_{\tau}\subseteq\overline{\mathcal{X}}_{\overline{\tau}}$.
Analogously, assume that there exists some constant $c_{\tau}\geq 0$ such that
$\min_{x,j,j^{\prime}}\Big{[}(\beta^{*}_{1j})^{\top}x-(\beta^{*}_{1j^{\prime}})^{\top}x\Big{]}=c^{*}_{\tau}\varepsilon,$
where the minimum is subject to $x\in\overline{\mathcal{X}}_{\overline{\tau}}$, $i\in\{\overline{\tau}_{1},\ldots,\overline{\tau}_{\overline{K}}\}$ and $i^{\prime}\in\{\overline{\tau}_{\overline{K}+1},\ldots,\overline{\tau}_{k}\}$. Then, if $M_{j}\leq\dfrac{c_{\tau}}{2}$, then we receive that $\overline{\mathcal{X}}_{\overline{\tau}}\subseteq\mathcal{X}^{*}_{\tau}$. As a consequence, by setting $M_{j}=\dfrac{1}{2}\min\{c^{*}_{\tau},c_{\tau}\}$, we achieve the conclusion that $\overline{\mathcal{X}}_{\overline{\tau}}=\mathcal{X}^{*}_{\tau}$. ∎
@@
$\sum_{i=1}^{k}\mathrm{softmax}(\mathrm{TopK}$ $(-\|W_{i}-X\|,K;\beta_{i}))\cdot f(Y|a_{i}^{\top}X+b_{i},\nu_{i})$
$=\sum_{i=1}^{k_{*}}\mathrm{softmax}(\mathrm{TopK}(-\|W^{*}_{i}-X\|,K;\beta^{*}_{i}))\cdot f(Y|(a^{*}_{i})^{\top}+b^{*}_{i},\nu^{*}_{i}).$ (32)
Due to the identifiability of the location-scale Gaussian mixtures [83, 84, 85], we get that $k=k_{*}$ and
$\Big{\{}\mathrm{softmax}(\mathrm{TopK}$ $(-\|W_{i}-X\|,K;\beta_{i})):i\in[k]\Big{\}}\equiv\Big{\{}\mathrm{softmax}(\mathrm{TopK}(-\|W^{*}_{i}-X\|,K;\beta^{*}_{i})):i\in[k]\Big{\}},$
@@
$\mathrm{softmax}(\mathrm{TopK}(-\|W_{i}-X\|,K;\beta_{i}))=\mathrm{softmax}(\mathrm{TopK}(-\|W^{*}_{i}-X\|,K;\beta^{*}_{i})),$ (33)
for almost surely $X$ for any $i\in[k]$. Since the $\mathrm{softmax}$ function is invariant to translations, it follows from equation 33 that $W_{i}=W^{*}_{i}$ and $\beta_{i}=\beta^{*}_{i}+v_{0}$ for some $v_{0}\in\mathbb{R}$. Notably, from the assumption of the model, we have $\beta_{0k}=\beta^{\prime}_{0k}=0$, which implies that $v_{0}=0$. As a result, we obtain that $\beta_{i}=\beta^{*}_{i}$ for any $i\in[k_{*}]$.
Let us consider $X\in\mathcal{X}_{\tau}$ where $\tau\in[q]$ such that $\{\tau_{1},\ldots,\tau_{K}\}=\{1,\ldots,K\}$. Then, equation L can be rewritten as
$\sum_{i=1}^{K}\exp(\beta_{i})\exp(-\|W_{i}-X\|)f(Y|a_{i}^{\top}X+b_{i},\nu_{i})$
$\hskip 85.35826pt=\sum_{i=1}^{K}\exp(\beta_{i})\exp(-\|W_{i}-X\|)f(Y|(a^{*}_{i})^{\top}X+b^{*}_{i},\nu^{*}_{i}),$ (34)
for almost surely $(X,Y)$. Next, we denote $J_{1},J_{2},\ldots,J_{m}$ as a partition of the index set $[k]$, where $m\leq k$, such that $\exp(\beta_{i})=\exp(\beta_{i^{\prime}})$ for any $i,i^{\prime}\in J_{j}$ and $j\in[m]$. On the other hand, when $i$ and $i^{\prime}$ do not belong to the same set $J_{j}$, we let $\exp(\beta_{i})\neq\exp(\beta_{i^{\prime}})$. Thus, we can reformulate equation L as
$\sum_{j=1}^{m}\sum_{i\in{J}_{j}}\exp(\beta_{i})\exp(-\|W_{i}-X\|)f(Y|a_{i}^{\top}X+b_{i},\nu_{i})$
$\hskip 85.35826pt=\sum_{j=1}^{m}\sum_{i\in{J}_{j}}\exp(\beta_{i})\exp(-\|W_{i}-X\|)f(Y|(a^{*}_{i})^{\top}X+b^{*}_{i},\nu^{*}_{i}),$
@@
$G=\sum_{j=1}^{m}\sum_{i\in J_{j}}\exp(\beta_{i})\delta_{(W_{i},a_{i},b_{i},\nu_{i})}=\sum_{j=1}^{m}\sum_{i\in J_{j}}\exp(\beta^{\prime}_{i})\delta_{(W^{*}_{i},a^{*}_{i},b^{*}_{i},\nu^{*}_{i})}=G_{*}.$

### v4

@@ I.3 Over-specified Setting
Arguing in a similar fashion as in Step 3 of Appendix K.1, we obtain that $\xi_{j,\alpha_{1},\alpha_{2},\eta}=\kappa_{j,\gamma}=0$ for any $j\in[k_{*}]$, $0\leq|\alpha_{1}|+|\alpha_{2}|\leq 2\bar{r}_{j}$, $0\leq\eta\leq 2(\bar{r}_{j}-|\alpha_{1}|-|\alpha_{2}|)$ and $0\leq|\gamma|\leq 1+\mathbf{1}_{\{|\mathcal{A}_{j}|>1\}}$. This contradicts the fact that at least one among them is non-zero. Hence, the proof is completed.
Appendix L Identifiability of the Laplace Gating Gaussian MoE
@@
$\sum_{i=1}^{k}\mathrm{softmax}($ $-\|W_{i}-X\|+\beta_{i})\cdot f(Y|a_{i}^{\top}X+b_{i},\nu_{i})$
$=\sum_{i=1}^{k_{*}}\mathrm{softmax}(-\|W^{*}_{i}-X\|+\beta^{*}_{i})\cdot f(Y|(a^{*}_{i})^{\top}+b^{*}_{i},\nu^{*}_{i}).$ (27)
Due to the identifiability of the location-scale Gaussian mixtures Teicher [1960, 1961, 1963], we get that $k=k_{*}$ and
$\Big{\{}\mathrm{softmax}(-\|W_{i}-X\|+\beta_{i}):i\in[k]\Big{\}}\equiv\Big{\{}\mathrm{softmax}(-\|W^{*}_{i}-X\|+\beta^{*}_{i}):i\in[k]\Big{\}},$
@@
$\mathrm{softmax}(-\|W_{i}-X\|+\beta_{i})=\mathrm{softmax}(-\|W^{*}_{i}-X\|+\beta^{*}_{i}),$ (28)
for almost surely $X$ for any $i\in[k]$. Since the $\mathrm{softmax}$ function is invariant to translations, it follows from equation 28 that $W_{i}=W^{*}_{i}$ and $\beta_{i}=\beta^{*}_{i}+v_{0}$ for some $v_{0}\in\mathbb{R}$. Notably, from the assumption of the model, we have $\beta_{k}=\beta^{*}_{k}=0$, which implies that $v_{0}=0$. As a result, we obtain that $\beta_{i}=\beta^{*}_{i}$ for any $i\in[k_{*}]$. Then, equation L can be rewritten as
$\sum_{i=1}^{k_{*}}\exp(\beta_{i})\exp(-\|W^{*}_{i}-X\|)f(Y|a_{i}^{\top}X+b_{i},\nu_{i})$
$\hskip 85.35826pt=\sum_{i=1}^{k_{*}}\exp(\beta_{i})\exp(-\|W^{*}_{i}-X\|)f(Y|(a^{*}_{i})^{\top}X+b^{*}_{i},\nu^{*}_{i}),$ (29)
for almost surely $(X,Y)$. Next, we denote $J_{1},J_{2},\ldots,J_{m}$ as a partition of the index set $[k_{*}]$, where $m\leq k_{*}$, such that $\exp(\beta_{i})=\exp(\beta_{i^{\prime}})$ for any $i,i^{\prime}\in J_{j}$ and $j\in[m]$. On the other hand, when $i$ and $i^{\prime}$ do not belong to the same set $J_{j}$, we let $\exp(\beta_{i})\neq\exp(\beta_{i^{\prime}})$. Thus, we can reformulate equation L as
$\sum_{j=1}^{m}\sum_{i\in{J}_{j}}\exp(\beta_{i})\exp(-\|W^{*}_{i}-X\|)f(Y|a_{i}^{\top}X+b_{i},\nu_{i})$
$\hskip 85.35826pt=\sum_{j=1}^{m}\sum_{i\in{J}_{j}}\exp(\beta_{i})\exp(-\|W^{*}_{i}-X\|)f(Y|(a^{*}_{i})^{\top}X+b^{*}_{i},\nu^{*}_{i}),$
@@
$G=\sum_{j=1}^{m}\sum_{i\in J_{j}}\exp(\beta_{i})\delta_{(W_{i},a_{i},b_{i},\nu_{i})}=\sum_{j=1}^{m}\sum_{i\in J_{j}}\exp(\beta^{*}_{i})\delta_{(W^{*}_{i},a^{*}_{i},b^{*}_{i},\nu^{*}_{i})}=G_{*}.$